OpenStudy (anonymous):

is this correct in simpify form / postive or negtive? -2/3 - [- 2/7] = 14/21 + 6/21= 20/21

OpenStudy (amistre64):

what is the [ ] notation represent here?

OpenStudy (anonymous):

-2/7

OpenStudy (amistre64):

so they are just brackets to indicate a grouping, ok .... remember this, 2 -s always is + -2/3 -- 2/7 -2/3 + 2/7 = some number -2(3)/3 + 2(3)/7 = (3)N -2 + 6/7 = 3N -2(7) + 6(7)/7 = 3(7)N -14 + 6 = 21N -12 = 21N -12/21 = N -4/7 = N is what i get

OpenStudy (anonymous):

thanks you

OpenStudy (amistre64):

yw